# Cavity-Altered Superconductivity

### v1

@@
Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment? Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [9, 10, 11, 12]. Here, we devised and implemented a novel platform to realize cavity-altered materials. Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and superior quality factor [13, 14, 15, 16, 17, 18, 19]. We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET). The frequencies of infrared (IR) hyperbolic modes of hBN [20, 21] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [22, 23]. Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET. Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface. Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the superfluid suppression. These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state. This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground-state properties of materials using IR-active phonons.
@@
Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice. The emergent effects born out of such strong interactions are abundant and drive the formation of a plethora of varied electronic and magnetic phases, as well as competing orders. At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter. A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring its quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8]. Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase of the electron phonon coupling via phonon polaritons [1, 18, 7], Amperean pairing directly mediated through dark-cavity photons [24, 25], and the driven-cavity extension of the Eliashberg effect [26]. Experiments with dark cavities have revealed altered ground states in the integer and fractional quantum Hall states of a 2D electron gas [9, 27, 28], as well as in charge density wave [10] systems, with additional reports suggesting changes to the diamagnetic response in superconducting nanoparticles [12, 11].
@@
We focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings. We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts with transition temperature () near 11.5 K [29, 30] (Supplementary Information Sections 1,2). We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN), interfaced with -ET (Fig. 1 bottom). Mode coupling at the interface is supported by the divergent photonic density of states [31, 32, 33] in the frequency range of hBN’s hyperbolicity (the permittivity has opposite signs along different axes [20, 21]), spanning the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [34]. Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode that is implicated in superconductivity [22, 23]. In this work, we examine how superconductivity is altered at the hBN/-ET interface, using scanning probe-based techniques to investigate local properties that are both inaccessible to bulk probes, and many surface probes due to the buried interfacial system.
@@
Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [36, 37, 38, 39]. The force results from the supercurrent screening the stray magnetic field of the tip, making MFM sensitive to the superfluid density near the superconductor surface, where “near” signifies the decay length for currents running parallel to the surface . Notably, the Meissner force is not affected by inserting nonmagnetic materials between the tip and the superconductor, and data in Fig. 1(a,b) attest that MFM probes superconductors encapsulated by insulating layers. We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [40, 41], and analysis of allows extraction of the superfluid density [38] (Supplementary Information Section 4).
We examine the impact of hBN on superconductivity in a heterostructure composed of a 60 nm thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick). We report measured as a function of , the separation between the tip and the -ET surface, in Fig. 1a. On a region of the -ET crystal un-obscured by hBN (green), at T=2 K, is negative, indicating a repulsive force. The magnitude of grows with reduced tip-sample surface separation. The measurement above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading. As a control experiment, we placed a 55 nm thick RuCl3 microcrystal on top of -ET. RuCl3 is an insulator with a static permittivity similar to hBN [42, 35]. However all optical phonons of RuCl3 occur at much lower frequencies below 350 cm-1 [35] and therefore are not expected to resonantly couple to the C=C stretching mode of -ET (Fig. 1c schematically depicts the permittivities of hBN, -ET, and RuCl3). Comparing measurements above RuCl3/-ET (Fig. 1b, magenta curve) and bare -ET (green curve - same as in a), we observe a much weaker (less than %) effect than over hBN/-ET. The stark contrast between the RuCl3/-ET and hBN/-ET interfaces suggests that the suppression of superfluid density under hBN is a result of frequency matching across the hBN/-ET interface. While hBN and RuCl3 may exert mechanical forces on the -ET surface, superfluid density has been shown to be minimally impacted under pressure [43]. We perform a second control measurement on a heterostructure of 8 nm hBN on a 28.5 nm thick microcrystal of (BSCCO) with K hosting phonons below 650 cm-1 [44], far below the hyperbolic modes in hBN (Supplementary Information Section 5). In Fig. 1b (inset), hBN/BSCCO (blue) is indistinguishable from bare BSCCO (dark gray). Thus, the significant suppression of the superfluid density is unique to the hBN/-ET interface.
@@
To experimentally examine the mode coupling at the hBN/-ET interface, we use scattering-type scanning near-field optical microscopy (s-SNOM) [48] to evaluate the hBN HPhP dispersion. We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser. The strong electric field with high momentum achieved near the tip apex launches polariton waves. We performed nano-IR measurements at K on a region of a hBN/-ET heterostructure with 32 nm hBN on 100 nm -ET. Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images, as shown in Fig. 2b for cm-1. Profiles through s-SNOM images collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig. 2c). In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with frequency [49]. However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the dashed line tracking the fifth peak in Fig 2c.
To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HPhP and the C=C stretching mode. First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP evanescent electric field with the C=C stretching mode. The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig 2d and Supplementary Information Section 2). Therefore, the C=C stretching mode does not effectively couple to the in-plane moments of the bare LO and TO phonons in hBN. In contrast, the HPhP evanescent field has a large out-of-plane component (Fig. 2d). To study the effect of this evanescent field, and its role in the renormalization of the zero-point motion of the C=C stretching mode, we calculated the C=C amplitude in the presence of an oscillating out-of-plane electric field induced by a phonon-polariton mode with a frequency spectrum (fundamental frequency cm-1), spanning the C=C frequency. With the temperature fixed at K, Langevin dynamics (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude, seen as a reduced amplitude and splitting of the peak in Fig. 2e. The negligible contribution of the direct phonon-phonon interaction is demonstrated by suppressing phonon-phonon scattering. Altogether, our Langevin molecular dynamics simulations indicate that cavity electrodynamics are crucial to explain the coupling between hBN and -ET. Importantly, vacuum fluctuations (zero-point phonon motion) and virtual processes invoked in mode hybridization permit the excitation of HPhPs without external photons, enabling coupling in the dark.
@@
We confirm the association of reduced (Fig 1a) with suppressed superfluid density by studying the temperature dependence. Over bare -ET, the large repulsive signal is well described by the Meissner effect (Supplementary Information Section 4), decreases with increasing temperature (Fig 3a) and vanishes near . Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effect at the heterostructure interfaces by performing constant height scans with the tip 300 nm above the -ET surface in a 1616 m2 region including both hBN and RuCl3 (Fig. 3b-d). To facilitate comparison across temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET. At 2 K (panel b), pN/m over hBN indicates a weaker Meissner force. Similar to the curves in Fig. 1(a,b), the effect over RuCl3 is negligible. Warming to 12 K, just above , the image contrast disappears, giving everywhere. The vanishing near confirms that the weakened observed at 2 K is associated with a reduction in the superfluid density. Further evaluation of measurement artifacts is available in Supplementary Information Section 8.
@@
We have studied multiple hBN/-ET devices that show a suppression of superfluid density under hBN. The effect varies spatially (Supplementary Information Section 9); we focus on regions with the strongest suppression. The results are summarized in Fig. 4. To model the Meissner force analytically [36], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth (schematic in Fig. 4 inset) and evaluate the effective superfluid density under hBN, . In Fig. 4, we compile the suppression observed across all devices as a function of the local hBN (or RuCl3) thickness, where is the superfluid density in bare -ET. The minimum value of extracted on each device (maximal suppression) is at least 50%, emphasizing the sizeable effect. This result holds for a large variation of the hBN thicknesses ranging from 25 nm to 110 nm. Accurate conversion between and superfluid density would require knowledge of the depth profile. While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate between possible depth profiles all leading to similar Meissner force. However, the magnitude of the observed Meissner force suppression implies that a significant thickness of -ET is impacted, i.e. a depth at least the same order of magnitude as . A lower bound on the depth of suppression is roughly 430 nm (Supplementary Information Section 10). In conclusion, our hBN/-ET structures reveal the suppression of superfluid density in the presence of resonant mode coupling across the interface.
Cavity electrodynamics
Our findings indicate that cavity electrodynamics plays a role in the suppression of superfluid density. Specifically: (1) The C=C stretching mode intersects the hBN HPhP dispersion in a regime of highly mixed light-like and phonon-like character. (2) Significant coupling to the C=C stretching mode dipole requires the out-of-plane evanescent field of the HPhP, which typically extends several hundred nanometers outside hBN. The poor screening of interlayer electric fields in -ET [34] therefore opens the possibility for coupling deep within the crystal. Consequently, it is plausible that dark cavity electrodynamics impact the superconducting ground state through coupling to the C=C stretching mode.
Our simulations indicate that hBN cavity modes reduce C=C stretching mode fluctuations, suggesting that enhancing fluctuations, e.g. as previously suggested through mode softening in the cavity [1], may increase in -ET. The feasibility of a cavity-coupled ground state in -ET highlights the potential of dark cavities in engineering quantum electronic ground states through the control of the quantum dynamics in the material via coupling to engineered cavity vacuum fluctuations [19, 50, 16, 18].
Although our results apply to the -ET superconducting ground state, the concept of employing tunable vdW hyperbolic cavities to resonantly couple materials to cavity modes is highly general. vdW hyperbolic cavities host modes across the IR and THz energy ranges [15], allowing resonant coupling to excitations across this range, and in turn, ground state engineering of many electronic orders.
@@
- [29] Ishiguro, T., Yamaji, K. & Saito, G. Organic superconductors (Springer-Verlag, Berlin, Heidelberg, 1998), 2nd edn.
- [30] Ardavan, A. et al. Recent Topics of Organic Superconductors. J. Phys. Soc. Jpn. 81, 11004 (2012).
@@
- [37] Roseman, M., Grütter, P., Badia, A. & Metlushko, V. Flux lattice imaging of a patterned niobium thin film. J. Appl. Phys. 89, 6787–6789 (2001).
@@
- [40] Schwarz, A. & Wiesendanger, R. Magnetic sensitive force microscopy. Nano Today 3, 28–39 (2008).
@@
- [43] Wakamatsu, K., Miyagawa, K. & Kanoda, K. Superfluid density versus transition temperature in a layered organic superconductor -(BEDT-TTF)2Cu[N(CN)2]Br under pressure. Phys. Rev. Res. 2, 043008 (2020).
@@
- [50] Le Dé, B., Eckhardt, C. J., Kennes, D. M. & Sentef, M. A. Cavity engineering of hubbard u via phonon polaritons. J.Phys. Mater. 5, 024006 (2022).

### v3

@@
Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment? Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8, 9], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [10, 11, 12, 13, 14, 15]. Here, we devised and implemented a novel platform to realize cavity-altered materials. Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and prominent mode confinement [16, 17, 18, 19, 20, 21, 22]. We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET). The frequencies of infrared (IR) hyperbolic modes of hBN [23, 24] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [25, 26]. Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm the presence of resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET. Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface. Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the superfluid suppression. These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state. This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground state properties.
@@
Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice. The emergent effects born out of such strong interactions are abundant and drive the formation of varied electronic and magnetic phases. At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter. A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring the quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8, 9]. Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase in electron-phonon coupling via phonon polaritons [1, 21, 7]; Amperean pairing directly mediated through dark-cavity photons [27, 28]; and the driven-cavity extension of the Eliashberg effect [29]. Experiments with dark cavities revealed modified transport properties in the integer and fractional quantum Hall states of a 2D electron gas [10, 13, 15] as well as cavity-assisted thermal control of the metal-to-insulator transition in charge density wave systems [11].
@@
We focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings. We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts below the transition temperature =11.5 K [30, 31] (Supplementary Information Sections 1,2). We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN). Hyperbolicity occurs when the permittivity has opposite signs along different axes [23, 24] and results in a highly enhanced photonic density of states [32, 33, 34]. Notably, hyperbolic modes of hBN overlap with the frequency of the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [35] and the two modes hybridize at the hBN/-ET interface. Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode of -ET that is implicated in superconductivity [25, 26]. In this work, we examine how superconductivity is altered at the hBN/-ET interface, using advanced scanning probe-based techniques that effectively analyze the electrodynamics of buried interfaces.
@@
Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [37, 38, 39]. The force results from the supercurrent screening the stray magnetic field of the tip. Notably, the Meissner force is not affected when non-magnetic materials are inserted between the tip and the superconductor, and the data in Fig. 1(a,b) attest that MFM probes superconductors encapsulated by insulating layers. We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [40]. Analysis of allows the extraction of the superfluid density [38] (Methods and Supplementary Information Section 4).
We examine the impact of hBN on superconductivity in a representative heterostructure composed of a 60-nm-thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick). We report measured as a function of , the separation between the tip and the -ET surface, in Fig. 1a. On a region of the -ET crystal unobscured by hBN (green), is negative at K, indicating a repulsive force. The magnitude of grows with reduced tip-sample surface separation. The signal collected above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading. As a control experiment, we placed a 55-nm-thick RuCl3 microcrystal on top of -ET. RuCl3 is an insulator with a static permittivity similar to that of hBN [41, 36]. However, optical phonons of RuCl3 all occur at much lower frequencies below 350 cm-1 [36] and therefore do not resonantly couple to the C=C stretching mode of -ET. Fig. 1c schematically depicts the permittivities of hBN, -ET, and RuCl3. Comparing measurements above RuCl3/-ET (Fig. 1b, magenta curve) and bare -ET (green curve - same as in Fig. 1a), we observe a much weaker (less than %) effect than over hBN/-ET. The stark contrast between the RuCl3/-ET and hBN/-ET interfaces indicates that the suppression of superfluid density under hBN is likely prompted by resonant coupling across the hBN/-ET interface. We remark that the RuCl3 control experiment rules out a prominent role for work function disparity [42] in the suppression of the superfluid density in Fig. 1 and similarly suggests that strain is unlikely to be significant. We perform a second control measurement on a heterostructure of 8-nm hBN residing on a 28.5-nm-thick microcrystal of (BSCCO) with K. Crystals of BSCCO host phonons below 650 cm-1 [43], far below the hyperbolic modes in hBN (Supplementary Information Section 5). MFM data plotted in Fig. 1b (inset) for hBN/BSCCO (blue) are indistinguishable from the data for bare BSCCO (dark gray). Thus, the significant suppression of the superfluid density is specific to the resonant nature of the hBN/-ET interface.
@@
To experimentally examine the mode coupling at the hBN/-ET interface, we used scattering-type scanning near-field optical microscopy (s-SNOM) [47]. Specifically, we evaluated the hBN HPhP dispersion as these modes couple to the C=C stretching mode of -ET. We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser. The strong electric field with high momentum achieved near the tip apex launches polariton waves. We performed nano-IR measurements at K on a region of an hBN/-ET heterostructure with 32-nm hBN on 100-nm -ET. Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images: representative results are shown in Fig. 2b for cm-1. Line profiles from s-SNOM collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig. 2c). In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with the frequency [48]. However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the solid line tracking the fifth peak in Fig. 2c. These nano-IR data attest to mode coupling across the hBN--ET interface.
To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HPhP and the C=C stretching mode. First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP’s electric field with the C=C stretching mode. The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig. 2d and Supplementary Information Section 2). It is important to note that the HPhP field has an out-of-plane component that can couple to the C=C mode (Fig. 2d). The combined contributions of many HPhP branches outspreading to high momenta make this field extend far beyond the hBN surface. We therefore calculate the C=C amplitude in the presence of an oscillating out-of-plane electric field with a frequency spectrum spanning the C=C frequency (fundamental frequency cm-1). Langevin molecular dynamics analysis carried out at K (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude and splits the peak in Fig. 2e. We therefore conclude that vacuum fluctuations—specifically, zero-point phonon motion—mediate the coupling between hyperbolic phonon polaritons and molecular vibrations without the need for external photons, thus enabling interaction in the dark.
@@
We now validate the association of reduced (Fig. 1a) with suppressed superfluid density by studying the temperature dependence. Over bare -ET, the large repulsive is rooted in the Meissner effect (Supplementary Information Section 4), it decreases with increasing temperature (Fig. 3a) and vanishes near . Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effects at the heterostructure interfaces by performing constant-height scans with the tip hovering above the -ET surface in a 1616-m2 region that includes both hBN and RuCl3 (Fig. 3b-d). To facilitate comparison between temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET. At 2 K (panel b), pN/m over hBN, indicating a weaker Meissner force. As with the curves in Fig. 1(a,b), the effect over RuCl3 is negligible. Warming to 12 K, just above , the image contrast disappears, giving everywhere. The vanishing near confirms that the weakened observed at 2 K is associated with a reduction in the superfluid density. Further evaluation and additional control measurements are described in Supplementary Information Section 8.
@@
We observe suppression of superfluid density under hBN in multiple hBN/-ET heterostructures. The magnitude of suppression varies spatially (Supplementary Information Section 9); we focus on regions with the strongest effect. The results are summarized in Fig. 4. To model the Meissner force analytically [37], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth and evaluate the effective superfluid density under hBN, . In Fig. 4, we compile MFM data for several devices with hBN cavities and RuCl3 dielectric layers of varied thicknesses. We present data in the form of , where is the superfluid density in bare -ET. In all heterostructures, we observe at least 50% suppression of the superfluid density prompted by the hBN cavity. This result holds over a large variation in the hBN thickness, ranging from 25 nm to 110 nm. Accurate conversion between and superfluid density would require knowledge of the depth profile (Supplementary Information Section 4). While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate among the possible depth profiles that all lead to similar Meissner forces. However, the magnitude of the observed Meissner force suppression implies that -ET is impacted by the cavity to a significant depth. The impacted depth appears to be of the same order of magnitude as the polaritonic wavelength (Supplementary Information Section 10).
Hyperbolic cavity electrodynamics
The data in Figs. 1-4 demonstrate that cavity electrodynamics plays a pivotal role in suppressing the superfluid density – a hallmark equilibrium property of superconductivity. These experiments establish a precedent for successfully altering a material’s ground state thermodynamic property through cavity engineering. Our specific implementation of the resonant structure reveals two notable features of hyperbolic cavity control. First, hyperbolic modes entail both in-plane and out-of-plane components of the fluctuating electromagnetic fields. The out-of-plane component is crucial for coupling to the out-of-plane dipole moment of the C=C stretching mode in the -ET superconductor under investigation. Second, the multiple branches of hyperbolic modes span a broad momentum range, enabling continuous interaction with the weakly dispersing C=C resonance across the entire phase space, as shown in Fig. 2a. Future experiments will be needed to evaluate the spatial extent of the mode coupling away from the hyperbolic interface; our earlier data for a similar resonant hBN/MoO3 heterostructure suggest that the relevant length scale may exceed nm [49].
Our data and modeling suggest a plausible pathway for cavity enhancement of superconducting properties. Indeed, our simulations indicate that the hBN cavity subdues the C=C stretching mode fluctuations prompting the stark suppression of superfluid density. By the same token, boosting fluctuations through mode softening in the cavity [1] may increase the superfluid density and possibly in -ET. Furthermore, the concept of employing readily reconfigurable vdW hyperbolic cavities for resonant manipulations of material properties is rather general. We remind the reader that vdW crystals host hyperbolic modes of phononic, plasmonic and exitonic origin [50] spanning the range from THz to visible, thus offering limitless options for hyper-cavity engineering.
@@
- [8] Flick, J., Rivera, N. & Narang, P. Strong light-matter coupling in quantum chemistry and quantum photonics. Nanophotonics 7, 1479–1501 (2018).
@@
- [30] Kanoda, K., Sakao, K., Takahashi, T., Komatsu, T. & Saito, G. NMR study of an organic superconductor, -(BEDT-TTF)2Cu [N(CN)2]Br. Physica C. 185, 2667–2668 (1991).
- [31] Dressel, M. et al. Electrodynamics of the organic superconductorsinebreak -(BEDT-TTF)2Cu(NCS)2 and -(BEDT-TTF)2Cu[N(CN)2]Br. Phys. Rev. B 50, 13603 (1994).
@@
- [42] Rizzo, D. J. et al. Charge-transfer plasmon polaritons at graphene/-RuCl3 interfaces. Nano lett. 20, 8438–8445 (2020).
@@
- [49] Sternbach, A. J. et al. Negative refraction in hyperbolic hetero-bicrystals. Science 379, 555–557 (2023).
- [50] Basov, D. et al. Polaritonic quantum matter. Nanophotonics (2025).